OpenStudy (noodya123):

A car rental company's standard charge includes an initial fee plus an additional fee for each mile driven. The standard charge S (in dollars) is given by the function S=17.75+0.40M, where is the number of miles driven. The company also offers an option to insure the car against damage. The insurance charge I (in dollars) is given by the function I=3.90+0.15m . Let C be the total charge (in dollars) for a rental that includes insurance. Write an equation relating C to M. Simplify your answer as much as possible. C=??

OpenStudy (luhvon):

did u get the answer yet ?

OpenStudy (noodya123):

no i didnt

OpenStudy (luhvon):

look in your messages

OpenStudy (noodya123):

never got a message from you

OpenStudy (welshfella):

S = 17.75 + 0.4M I = 3.90 + 0.15M C = S + I

OpenStudy (welshfella):

so to find C in terms of M you need to add the above 2 functions.

OpenStudy (luhvon):

I just sent it to u again

OpenStudy (noodya123):

so would it be 21.65 + 0.55M

OpenStudy (noodya123):

C=21.65=0.55M?

OpenStudy (welshfella):

yes

OpenStudy (noodya123):

thank you guys it was correct

OpenStudy (welshfella):

C = 21.65 + 0.55M

OpenStudy (welshfella):

yw